# Hessian stability and convergence rates for entropic and Sinkhorn potentials via semiconcavity

### v1

@@ 1.1. Exponential convergence of Hessian of Sinkhorn’s iterates
As already mentioned in the introduction, most of the popularity EOT has recently gotten is due to the possibility of rapidly computing its solutions via an iterative algorithm, known as Sinkhorn’s algorithm [55, 56] or Iterative Proportional Fitting Procedure. Given any initialization $\varphi^{0}\colon\mathbb{R}^{d}\to\mathbb{R}$, this algorithm solves (1.1) as a fixed point problem by generating two sequences $\{\varphi^{n},\psi^{n}\}_{n\in\mathbb{N}}$, called Sinkhorn potentials, defined recursively as:
$\begin{cases}\varphi^{n+1}=U_{\rho}+\log P_{T}\exp(-\psi^{n})\\ \psi^{n+1}=U_{\mu}+\log P_{T}\exp(-\varphi^{n+1})\;.\end{cases}$
@@
As pointed out in [5], this is also equivalent to the Bregman’s iterated projection algorithm for relative entropy. Indeed, in the current setup Bregman’s iterated projection algorithm produces two sequences of plans $(\pi^{n,n},\pi^{n+1,n})_{n\in\mathbb{N}}$ starting from a positive measure $\pi^{0,0}$ according to the following recursion:
$\pi^{n+1,n}\coloneqq{\arg\min}_{\Pi(\mu,\star)}\mathscr{H}(\cdot|\pi^{n,n})\,,\qquad\pi^{n+1,n+1}\coloneqq{\arg\min}_{\Pi(\star,\nu)}\mathscr{H}(\cdot|\pi^{n+1,n})\,,$
@@
where $\Pi(\rho,\star)$ (resp. $\Pi(\star,\mu)$) is the set of probability measures $\pi$ on $\mathbb{R}^{2d}$ such that the first marginal is $\rho$, i.e., $(\mathrm{proj}_{x})_{\sharp}\pi=\rho$ (resp. the second marginal is $\mu$, i.e., $(\mathrm{proj}_{y})_{\sharp}\pi=\mu$). It is relatively easy (cf. [45, Section 6]) to show that, starting from $\pi^{0,0}(\mathrm{d}x\mathrm{d}y)\propto\exp(-\nicefrac{{|x-y|^{2}}}{{2T}}-\psi^{0}(y)-\varphi^{0}(x))\mathrm{d}x\mathrm{d}y$, the iterates in (1.1) are related to Sinkhorn potentials through
$\textstyle\pi^{n+1,n}(\mathrm{d}x\mathrm{d}y)$ $\textstyle\propto\exp(-\nicefrac{{|x-y|^{2}}}{{2T}}-\varphi^{n+1}(x)-\psi^{n}(y))\mathrm{d}x\mathrm{d}y\;,$
$\textstyle\pi^{n+1,n+1}(\mathrm{d}x\mathrm{d}y)$ $\textstyle\propto\exp(-\nicefrac{{|x-y|^{2}}}{{2T}}-\varphi^{n+1}(x)-\psi^{n+1}(y))\mathrm{d}x\mathrm{d}y\;.$
@@
In the sequel, we will refer to the couplings $(\pi^{n,n},\pi^{n+1,n})_{n\in\mathbb{N}}$ as Sinkhorn plans. By definition $\pi^{n+1,n}$ has the correct first marginal, but wrong second marginal, which we denote with $\mu^{n+1,n}$. Similarly, the second marginal of $\pi^{n,n}$ is fitted, however the first one might not be correct and hereafter we will denote it as $\rho^{n,n}$. Moreover, $\pi^{n+1,n}$ is the optimal entropic plan associated to the EOT problem with marginals $\rho,\,\mu^{n+1,n}$ whereas $\pi^{n,n}$ is the optimal EOT associated to the problem with marginals $\rho^{n,n},\,\mu$. Due to this partial marginal fitting nature of the algorithm, since we can see Sinkhorn plans $\{\pi^{n+1,n}\}_{n\in\mathbb{N}}$ as a sequence of entropic plans where the first marginal is always fixed and the second one changes according to $\{\mu^{n+1,n}\}$, we see that proving the exponential convergence of the algorithm boils down to apply quantitative stability estimates and to control the sequence of wrong marginals. For these reasons, Sinkhorn’s algorithm and quantitative convergence bounds for EOT are two problems tightly related and both have been addressed from a vast literature (see literature review below). Despite this, in the unbounded settings, much less has been known until the recent contributions of [15, 28, 13], where this problem has been addressed in full generality and where exponential convergence rates were shown to hold in relative entropy for Sinkhorn plans and in $\mathrm{L}^{p}$-norm (with $p\in\{1,2\}$) for gradients of Sinkhorn potentials. Here our Hessian stability estimates allow us to deduce also a second order convergence result, i.e., that the Hessian of Sinkhorn potentials converges exponentially fast with the same rate obtained in [13] for Sinkhorn plans. To state it, let us recall that a probability measure $\nu\in\mathcal{P}(\mathbb{R}^{d})$ is said to satisfy a Talagrand’s inequality with constant $\tau$, TI($\tau$) for short, if
@@
Assume H 1 and that there exist $\Lambda\in(0,+\infty)$ and $N\geq 2$ such that
$y\mapsto g^{x}_{\psi_{0}^{n}}(y)=\frac{|x-y|^{2}}{2}-T\,\varphi_{0}^{n}(y)$
is $\Lambda$-semiconcave [^2] This is the same assumption considered in [13] when proving the exponential convergence of Sinkhorn’s plans. Here we have a $-T$ in front of $\psi_{0}$ due to notational difference. uniformly in $x\in\operatorname{supp}(\rho)$ and $n\geq N$. If $\mu^{n,n-1}$ satisfies TI($\tau$) for some $\tau\in(0,+\infty)$ and for all $n\geq N$, then
@@
$\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\lesssim\tau\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,,$
@@
$\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\lesssim\sqrt{\tau}\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{\frac{n-N+1}{2}}\,\sqrt{\mathscr{H}(\pi^{\mu}|\pi^{0,0})}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\tau\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,.$
@@
hold for all $n\geq N$ up to multiplicative constants that depend polynomially only on $\rho,\,\mu,\,T$ (and not on the iterates). These constants are explicit and in particular, up to numerical universal constants, we have
@@
  • if $\operatorname{supp}(\rho),\,\operatorname{supp}(\mu)\subseteq B_{R}(0)$ (for some radius big enough, i.e., $R^{2}\geq T$) then
@@
    $\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\lesssim\tau\,\nicefrac{{R^{4}}}{{T^{4}}}\,\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,,$
@@
    $\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\lesssim\sqrt{\tau}(\nicefrac{{R^{4}}}{{T^{\nicefrac{{7}}{{2}}}}}+\nicefrac{{d}}{{T}})\,\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{\frac{n-N+1}{2}}\,\sqrt{\mathscr{H}(\pi^{\mu}|\pi^{0,0})}$
    $\qquad\qquad\qquad\qquad\qquad\qquad+\tau\,\nicefrac{{R^{6}}}{{T^{5}}}\,\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,,$
@@
  • if both $\rho$ and $\mu$ are log-concave, i.e., their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\mu}\geq\alpha_{\mu}$ for some $\alpha_{\rho},\,\alpha_{\mu}>0$ (wlog such that $\alpha_{\rho}\vee\alpha_{\mu}<T^{-1}$), then
@@
    $\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\lesssim\frac{\tau}{\alpha_{\rho}\,\alpha_{\mu}\,T^{4}}\,\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,,$
@@
    $\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\lesssim\sqrt{\tau}\biggl{(}\frac{1}{\alpha_{\mu}\,\sqrt{\alpha_{\rho}}\,T^{3}}+\frac{d}{\sqrt{\alpha_{\rho}\,\alpha_{\mu}}\,T^{2}}\biggr{)}\,\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{\frac{n-N+1}{2}}\,\sqrt{\mathscr{H}(\pi^{\mu}|\pi^{0,0})}$
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{\tau}{\alpha_{\rho}\,\alpha_{\mu}\,T^{4}}\,\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,.$
Lastly, let us remark that the multiplicative constants appearing in this theorem are the same ones obtained in Theorem 1.1 (exchanging the role between $\nu$ and $\mu$).
1.2. The Schrödinger bridges point of view
@@
Our proof strategy relies on the stochastic control representation of entropic plans as laws of solutions to time-inhomogeneous SDEs. More precisely, we are going to consider the forward Schrödinger bridge process (from $\rho$ to $\mu$) defined as the SDE driven by $-\nabla\psi^{\mu}_{s}$, that is the stochastic process $(X^{\psi^{\mu},\rho})_{s\in[0,T]}$ solution to (cf. [20])
@@
Similarly, we will consider its time reversal corresponding process, i.e., the (backward) Schödinger bridge (from $\mu$ to $\rho$) which solves
@@
In light of these representation, it is clear that semiconcavity and functional properties of the EOT plan $\pi^{\mu}$ are affected by convexity properties of the drifts appearing in (1.3) and (1.4), as already noticed in [20]. For this reason, alongside the semiconcavity parameter $\Lambda(\varphi^{\mu}_{0})$ our constants appearing below will depend on lower bounds on the Hessians of propagated potentials along HJB, i.e., for any $h\in\{\varphi^{\mu},\,\psi^{\mu}\}$ we will consider the lower bounds $\nabla^{2}h_{s}\geq\lambda(h_{s})$ with $\lambda(h_{s})\in\mathbb{R}$. Our general results are stated for any given sequence $(\lambda(h_{s}))_{s\in[0,T)}$ satisfying this lower bound (and we do not assume it to be the optimal one, as we did for $\Lambda(\varphi^{\mu}_{0})$). In Appendix A and Appendix B we will provide explicit lower bounds for the examples considered here.

### v2

@@ 1.1. Exponential convergence of Hessian of Sinkhorn’s iterates
Most of the popularity EOT has recently gotten is due to the possibility of rapidly computing its solutions via an iterative algorithm, known as Sinkhorn’s algorithm [Sin64, SK67] or Iterative Proportional Fitting Procedure (IPFP). Given any initialization $\varphi^{0}\colon\mathbb{R}^{d}\to\mathbb{R}$, this algorithm solves (1.1) as a fixed point problem by generating two sequences $\{\varphi^{n},\psi^{n}\}_{n\in\mathbb{N}}$, called Sinkhorn potentials, defined recursively as:
$\varphi^{n+1}=U_{\rho}+\log P_{T}\exp(-\psi^{n})\,,\qquad\psi^{n+1}=U_{\mu}+\log P_{T}\exp(-\varphi^{n+1})\;.$
@@
As pointed out in [BCC+15], this is also equivalent to Bregman’s iterated projection algorithm for relative entropy, which in the current setup produces two sequences of plans $(\pi^{n,n},\pi^{n+1,n})_{n\in\mathbb{N}}$ starting from a positive measure $\pi^{0,0}$ according to the following recursion:
$\pi^{n+1,n}\coloneqq{\arg\min}_{\Pi(\rho,\star)}\mathscr{H}(\cdot|\pi^{n,n})\,,\qquad\pi^{n+1,n+1}\coloneqq{\arg\min}_{\Pi(\star,\mu)}\mathscr{H}(\cdot|\pi^{n+1,n})\,$,
@@
where $\Pi(\rho,\star)$ (resp. $\Pi(\star,\mu)$) is the set of probability measures $\pi$ on $\mathbb{R}^{2d}$ such that the first marginal is $\rho$, i.e., $(\mathrm{proj}_{x})_{\sharp}\pi=\rho$ (resp. the second marginal is $\mu$, i.e., $(\mathrm{proj}_{y})_{\sharp}\pi=\mu$). It is relatively easy (cf. [Nut21, Section 6]) to show that, starting from $\pi^{0,0}(\mathrm{d}x\mathrm{d}y)\propto\exp(-\nicefrac{{|x-y|^{2}}}{{2T}}-\psi^{0}(y)-\varphi^{0}(x))\mathrm{d}x\mathrm{d}y$, the iterates in (1.1) are related to Sinkhorn potentials through
$\textstyle\pi^{n+1,n}(\mathrm{d}x\mathrm{d}y)\textstyle\propto e^{-\frac{|x-y|^{2}}{2T}-\varphi^{n+1}(x)-\psi^{n}(y)}\mathrm{d}x\mathrm{d}y\;,\qquad\textstyle\pi^{n+1,n+1}(\mathrm{d}x\mathrm{d}y)\textstyle\propto e^{-\frac{|x-y|^{2}}{2T}-\varphi^{n+1}(x)-\psi^{n+1}(y)}\mathrm{d}x\mathrm{d}y\;.$
@@
In the sequel, we will refer to the couplings $(\pi^{n,n},\pi^{n+1,n})_{n\in\mathbb{N}}$ as Sinkhorn plans. By definition $\pi^{n+1,n}$ has the correct first marginal, but wrong second marginal, which we denote with $\mu^{n+1,n}$. Similarly, the second marginal of $\pi^{n,n}$ is fitted, however the first one might not be correct and hereafter we will denote it as $\rho^{n,n}$. Moreover, $\pi^{n+1,n}$ is the optimal entropic plan associated to the EOT problem with marginals $\rho,\,\mu^{n+1,n}$ whereas $\pi^{n,n}$ is the optimal EOT plan associated to the problem with marginals $\rho^{n,n},\,\mu$. Due to this partial marginal fitting nature of the algorithm, since we can see Sinkhorn plans $\{\pi^{n+1,n}\}_{n\in\mathbb{N}}$ as a sequence of entropic plans where the first marginal is always fixed and the second one changes according to $\{\mu^{n+1,n}\}$, we see that proving the exponential convergence of the algorithm boils down to apply quantitative stability estimates and to control the sequence of wrong marginals. For these reasons, Sinkhorn’s algorithm and quantitative convergence bounds quantitative stability bounds for EOT are two problems tightly related and both have been addressed from a vast literature (see literature review below). Despite this, in the unbounded settings, much less has been known until the recent contributions of [CDG23, Eck25, CCGT24], where this problem has been addressed in full generality and where exponential convergence rates were shown to hold in relative entropy for Sinkhorn plans and in $\mathrm{L}^{p}$-norm (with $p\in\{1,2\}$) for gradients of Sinkhorn potentials. Here our Hessian stability estimates allow us to deduce also a second-order convergence result, i.e., that the Hessian of Sinkhorn potentials converges exponentially fast with the same rate obtained in [CCGT24] for Sinkhorn plans. To state it, let us recall that a probability measure $\nu\in\mathcal{P}(\mathbb{R}^{d})$ is said to satisfy a Talagrand inequality with constant $\tau$, TI($\tau$) for short, if
@@
Assume H 1 and that there exist $\Lambda\in(0,+\infty)$ and $N\geq 2$ such that the function $g^{x}_{\varphi_{0}^{n}}$ is $\Lambda$-semiconcave uniformly in $x\in\operatorname{supp}(\rho)$ and $n\geq N$. If $\mu^{n,n-1}$ satisfies TI($\tau$) for some $\tau\in(0,+\infty)$ and for all $n\geq N$, then
@@
$\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\lesssim\bigg(1-\frac{T}{T+\tau\Lambda}\bigg)^{(n-N+1)}\tau\,\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,$,
@@
$\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\lesssim\bigg(1-\frac{T}{T+\tau\Lambda}\bigg)^{\frac{n-N+1}{2}}\,\sqrt{\tau\mathscr{H}(\pi^{\mu}|\pi^{0,0})}+\bigg(1-\frac{T}{T+\tau\Lambda}\bigg)^{(n-N+1)}\tau\mathscr{H}(\pi^{\mu}|\pi^{0,0}),$
@@
hold for all $n\geq N$ up to multiplicative constants that depend polynomially only on $\rho,\,\mu,\,T$ (and not on the iterates). These constants are explicit.
In particular, up to numerical universal constants, we have
@@
  • if $\operatorname{supp}(\rho),\,\operatorname{supp}(\mu)\subseteq B_{R}(0)$ (for some radius big enough, i.e., $R^{2}\geq T$), then the uniform semiconcavity parameter reads as $\Lambda=R^{2}\,T^{-1}$ and
@@
    $\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\lesssim\nicefrac{{R^{4}}}{{T^{4}}}\,\bigg(1-\frac{T^{2}}{T^{2}+\tau R^{2}}\bigg)^{(n-N+1)}\tau\,\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,$,
@@
    $\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\lesssim(\nicefrac{{R^{4}}}{{T^{\nicefrac{{7}}{{2}}}}}+\nicefrac{{d}}{{T}})\,\bigg(1-\frac{T^{2}}{T^{2}+\tau R^{2}}\bigg)^{\frac{n-N+1}{2}}\,\sqrt{\tau\mathscr{H}(\pi^{\mu}|\pi^{0,0})}$
    $\qquad\qquad\qquad\qquad\qquad\qquad+\nicefrac{{R^{6}}}{{T^{5}}}\,\bigg(1-\frac{T^{2}}{T^{2}+\tau R^{2}}\bigg)^{(n-N+1)}\tau\,\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,$,
@@
  • if both $\rho$ and $\mu$ are log-concave, i.e., their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\mu}\geq\alpha_{\mu}$ for some $\alpha_{\rho},\,\alpha_{\mu}>0$ (wlog such that $\alpha_{\rho}\vee\alpha_{\mu}<T^{-1}$), then the uniform semiconcavity parameter reads as $\Lambda=(\alpha_{\rho}\,T)^{-1}$ and
@@
    $\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\lesssim\frac{\tau}{\alpha_{\rho}\,\alpha_{\mu}\,T^{4}}\,\bigg(1-\frac{\alpha_{\rho}\,T^{2}}{\alpha_{\rho}\,T^{2}+\tau}\bigg)^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,$,
@@
    $\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\lesssim\biggl(\frac{1}{\alpha_{\mu}\,\sqrt{\alpha_{\rho}}\,T^{3}}+\frac{d}{\sqrt{\alpha_{\rho}\,\alpha_{\mu}}\,T^{2}}\biggr)\,\bigg(1-\frac{\alpha_{\rho}\,T^{2}}{\alpha_{\rho}\,T^{2}+\tau}\bigg)^{\frac{n-N+1}{2}}\,\sqrt{\tau\mathscr{H}(\pi^{\mu}|\pi^{0,0})}$
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{\tau}{\alpha_{\rho}\,\alpha_{\mu}\,T^{4}}\,\bigg(1-\frac{\alpha_{\rho}\,T^{2}}{\alpha_{\rho}\,T^{2}+\tau}\bigg)^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,.$
Let us remark that the uniform $\Lambda$-semiconcavity of $g^{x}_{\varphi^{n}_{0}}$ (as defined in (1.3)) and the Talagrand inequality are the same assumptions considered in [CCGT24] when proving the exponential convergence of Sinkhorn’s plans, and combining them with our stability bounds for gradients and Hessians leads to the above first- and second-order convergence for Sinkhorn’s iterates. Moreover, as shown in Section A.1 and Section A.2, whenever $\rho$ is either compactly supported or log-concave, we are guaranteed that the function $g^{x}_{\varphi^{n}_{0}}$ is $\Lambda$-semiconcave uniformly in $n\in\mathbb{N}$, whence the validity of the first assumption in the previous theorem.
As concerns the uniform Talagrand inequality assumption on the marginals $\mu^{n,n-1}$ generated along Sinkhorn’s algorithm, let us comment its role and a possible alternative assumption. As for the role, we employ Talagrand inequality for a twofold reason: to apply [CCGT24] and to translate their entropic convergence estimates into our $\mathbf{W}_{2}$-stability bounds (cf. (3.6)). As for the alternative assumption, our convergence result can still be obtained by alternatively assuming $\mu$ to satisfy a Talagrand inequality (in order to apply [CCGT24, Theorem 1.2]) and Sinkhorn’s plans $\pi^{n,n-1}$ to satisfy a Talagrand inequality (in order to translate entropic bounds into Wasserstein ones). This last condition is met if for instance $\rho$ satisfies a log-Sobolev inequality, which implies its validity also for Sinkhorn’s plans $\pi^{n,n-1}$ [Con24, Theorem 1.3], and hence a Talagrand inequality (see for instance [BGL13, Theorem 9.6.1]).
1.2. Schrödinger bridge point of view
@@
Our proof strategy relies on the stochastic control representation of entropic plans as laws of solutions to time-inhomogeneous SDEs. More precisely, we are going to consider the forward Schrödinger bridge process (from $\rho$ to $\mu$) defined as the SDE driven by $\nabla\psi^{\mu}_{s}$, that is the stochastic process $(X^{\psi^{\mu},\rho})_{s\in[0,T]}$ solution to (cf. [Con24])
@@
Similarly, we will consider its time-reversal corresponding process, i.e., the (backward) Schrödinger bridge (from $\mu$ to $\rho$) which solves
@@
In light of these representations, it is clear that semiconcavity and functional properties of the EOT plan $\pi^{\mu}$ are affected by convexity properties of the drifts appearing in (1.4) and (1.5), as already noticed in [Con24]. For this reason, alongside the semiconcavity parameter $\Lambda(\varphi^{\mu}_{0})$ our constants appearing below will depend on lower bounds on the Hessians of propagated potentials along HJB, i.e., for any $h\in\{\varphi^{\mu},\,\psi^{\mu}\}$ we will consider the lower bounds $\nabla^{2}h_{s}\geq\lambda(h_{s})$ with $\lambda(h_{s})\in\mathbb{R}$. Our general results are stated for any given sequence $(\lambda(h_{s}))_{s\in[0,T)}$ satisfying this lower bound (and we do not assume it to be the optimal one, as we did for $\Lambda(\varphi^{\mu}_{0})$). In the Appendix we provide explicit lower bounds for the examples considered here.